whats three times the larger of two numbers is equal to four times the smaller. The sum of the numbers is 21. Find the number. Please show the work.
Answer by checkley77(12569)

X+Y=21 OR X=21-Y
3X=4Y
3(21-Y)=4Y
63-3Y=4Y
-3Y-4Y=-63
-7Y=-63
Y=-63/-7
Y=9 ANS.
X=21-9=12 ANS.
PROOF:
3*12=4*9
36=36